OpenStudy (anonymous):

A cylindrical vase has a diameter of 4 inches. At the bottom of the vase, there are 6 marbles, each of diameter 3 inches. The vase is filled with water up to a height of 8 inches. Which of the following could be used to calculate the volume of water in the vase? π(2in)2(8in) − 6(four over threeπ(1.5in)3) π(8in)2(2in) − 6(four over threeπ(1.5in)3) π(2in)2(8in) − 1.5(four over threeπ(6in)3) π(8in)2(2in) − 1.5(four over threeπ(6in)3)

OpenStudy (anonymous):

@geekfromthefutur

OpenStudy (anonymous):

@robtobey can u please help me ??

OpenStudy (anonymous):

The volume of a the cylinder minus the volume of 6 each, 3 inch diameter spheres.

OpenStudy (anonymous):

which is the first answer

OpenStudy (anonymous):

um.. i think it would be C im not sure though ?

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

V=4/3 * π* r^3

OpenStudy (anonymous):

volume of sphere

OpenStudy (anonymous):

A.\[\pi 2^2 8-\frac{6}{3} \left(4 \pi 1.5^3\right) \]LaTex is not rendering the input to it, correctly.

OpenStudy (anonymous):

oh ok thank yo guys so very much ^w^ could yall help me with just one more Q :D

OpenStudy (anonymous):

you*

OpenStudy (anonymous):

u should be giving medals

OpenStudy (anonymous):

oh sorry lol i almost forgot :p